Is angular acceleration about any point on a rotating rigid body same? If so what is the value of it for a rod of length with a force of M*g acting on the centre and M*g/2 acting on one end (both forces are in opposite direction)?

Grade:12

2 Answers

Vikas TU
7 years ago
net moment = I * a where a is the angular accln.
Therefore, moment about centre = Mg*l/2 = Ml^2/12 * a
Get the value of a Hence.
Guru Prasath
7 years ago
You dint get my question.. Think about the first part of my question.. If the answer is yes then there is a confusion for me..